OpenStudy (anonymous):

The number of users of the internet in a town increased by a factor of 1.01 every year from 2000 to 2010. The function below shows the number of internet users f(x) after x years from the year 2000: f(x) = 3000(1.01)^x Which of the following is a reasonable domain for the function? 0 ≤ x ≤ 10 2000 ≤ x ≤ 2010 0 ≤ x ≤ 3000 All positive integers

OpenStudy (jameshorton):

is there like a A or b @TheDragon

OpenStudy (anonymous):

Yeah. The first one is a, the second is b, the third is c, and the fourth d

OpenStudy (jameshorton):

ok A

OpenStudy (anonymous):

huh. okay can you help me on a second one?

OpenStudy (jameshorton):

ya

OpenStudy (anonymous):

A store had 175 cell phones in the month of January. Every month, 10% of the cell phones were sold and 10 new cell phones were stocked in the store. Which recursive function best represents the number of cell phones in the store f(n) after n months?

OpenStudy (anonymous):

f(n) = 175 - 0.9 × f(n - 1) + 10, f(0) = 175, n > 0 f(n) = 0.1 × f(n - 1) + 10, f(0) = 175, n > 0 f(n) = 175 + 0.9 × f(n - 1) + 10, f(0) = 175, n > 0 f(n) = 0.9 × f(n - 1) + 10, f(0) = 175, n > 0

OpenStudy (jameshorton):

D

OpenStudy (anonymous):

wait how?